# Trustworthy Enhanced Multi-view Multi-modal Alzheimer’s Disease Prediction with Brain-wide Imaging Transcriptomics Data

### v1

@@ III-C Modality-level Confidence: True-False-Harmonized Class Probability
In multimodal learning, data heterogeneity is a common challenge, with the discriminative capability of each modality varying across different samples. Therefore, it is crucial to develop algorithms that can adaptively perceive and respond to these variations in informativeness. By transforming the assessment of modal informativeness into an evaluation of the confidence levels associated with modal classification performance, we can effectively estimate the modality informativeness through its prediction confidence. Here, we propose a novel TFCP criterion to approximate the prediction confidence of each modality.
The confidence criterion measures how confident the model is in its predictions by correlating high prediction certainty with greater values and vice versa. For a classifier $f:\mathbf{x}_{i}\rightarrow y_{i}$, the traditional confidence criterion utilizes the maximum class probability (MCP): $\text{MCP}(\mathbf{x}_{i})=P(\widehat{y_{i}}|\mathbf{x}_{i})$, where $\widehat{y_{i}}$ is the class with the largest softmax probability. However, it can be observed that this method assigns high confidence (i.e., high softmax probability) to both correct and incorrect predictions, resulting in overconfidence for failure predictions. True class probability (TCP) criterion [37] has been proposed to calibrate incorrect predictions. Instead of assigning the maximum class probability, TCP uses the softmax probability of the true class $y_{i}^{*}$ as prediction confidence: $\text{TCP}(\mathbf{x}_{i})=P(y_{i}^{*}|\mathbf{x}_{i})$, where $y_{i}^{*}$ is the true class (i.e., the actual label). The TCP strategy exclusively models the certainty of the true class while neglecting the uncertainty of the untrue classes, which could result in a biased and unstable confidence approximation.
To this end, we propose considering both true and false class probabilities to aggregate evidence from these two perspectives. Specifically, we design a harmonized criterion for approximating the prediction confidence:
$\text{TFCP}(\mathbf{x}_{i})$ $=\frac{2}{1/{\text{TCP}(\mathbf{x}_{i})}+1/({1-\text{FCP}(\mathbf{x}_{i})})}$ (5)
@@
where $\text{FCP}(\mathbf{x}_{i})=P(\overline{y_{i}^{*}}|\mathbf{x}_{i})$ is the softmax probability of false (or untrue) class.
It should be noted that neither the TCP nor the FCP can be directly applied to test samples because they both require label information. To address this issue, we introduce two confidence networks to approximate the TCP and FCP, respectively. As shown in Figure 1(e), a TCP confidence network with parameters $\theta_{T}$ and an FCP confidence network with parameters $\theta_{F}$ are constructed based on the sample representations to generate certainty and uncertainty scores. These networks are trained to minimize the discrepancy between predicted and actual scores:
$\mathcal{L}_{\mathrm{Conf}}^{m}\!=\!\left\|\text{TFCP}(\mathbf{Z}^{m})-\widehat{\text{TFCP}}(\mathbf{Z}^{m},\theta^{m}_{T},\theta^{m}_{F})\right\|^{2}\!+\!\mathcal{L}_{\text{Cls}}^{m}.$ (6)
Here, both the TCP and FCP networks are built upon a classification network trained using the cross-entropy loss $\mathcal{L}_{\text{Cls}}^{m}$.
@@ IV-A Dataset
AHBA dataset. Brain-wide transcriptomics data are sourced from the Allen human brain atlas (AHBA, human.brain-map.org), including over 58k probes sampled across 3,702 brain locations from six donors. We use the abagen toolbox [39] for data preprocessing and employ the AAL atlas to map brain locations to ROIs, resulting in a number of 15,633 genes expressed across 116 ROIs. To ensure that the prior knowledge introduced closely aligns with the underlying mechanisms of AD, we further filter genes based on a large-scale AD meta-GWAS from the IGAP [38]. Specifically, we use the MAGMA [40] to derive gene-level $p$-values from SNP-level GWAS and keep nominally significant ones (i.e., $p<$0.05). This process yields a total of 1,216 genes across 116 ROIs for constructing the T-RRI edge matrix.
@@
NC: normal control; EMCI: early mild cognitive impairment; LMCI: late mild cognitive impairment; AD: Alzheimer’s disease.
ADNI dataset. We conduct experimental validation using subjects from the Alzheimer’s Disease Neuroimaging Initiative (ADNI, https://adni.loni.usc.edu/). Multimodal imaging data–including AV45-PET, FDG-PET, and VBM-MRI, along with corresponding diagnosis labels (NC, EMCI, LMCI, and AD) for each visit, are collected from 851 ADNI participants. Each scan is processed according to established protocols [4], and measurements for 116 ROIs are derived from each modality using the AAL atlas. Detailed characteristic information is listed in Table I.
Benchmark methods. We compare TMM with nine multimodal competitors, including three single-modal classifiers with early fusion (SVM, Lasso, XGBoost, and fully connected NN), three models with intermediate concatenation (GRidge [41], BSPLSDA [42], and GMU [43]), and two methods with advanced representation and fusion designs (Mogonet [44] and Dynamics [20]).
@@
Best results are in Bold. Suboptimal results are underlined. ‘*’ indicates that our TMM is significantly better ($p<0.05$) than the suboptimal method when using the independent t-test.
@@
Evaluation metrics. We employ accuracy (ACC), F1 score (F1), and the area under the receiver operating characteristic curve (AUC) to evaluate the performance of the methods. Five-fold cross-validation is performed to calculate the mean and standard deviation of the results. T-test is performed to evaluate the significance of improvements achieved by our method over state-of-the-art methods.
Implementation details. We develop the model using PyTorch 2.1.0 and Adam as the optimizer. We train the model for 2,000 epochs with a learning rate of 1e-3 and weight decay of 1e-4. All experiments are implemented on an RTX 4090 GPU with 24GB of memory.
IV-B Comparison with the state-of-the-art
Four tasks, including NC vs. AD, NC vs. LMCI, NC vs. EMCI, and EMCI vs. LMCI, are designed to validate the performance of our proposed method. Since the transcriptomics knowledge is not specific to ADNI subjects and cannot be directly incorporated, we are limited to using only multimodal imaging data for the comparative methods. Table II shows the comparison results where our model illustrates superior performance across four tasks. Particularly, TMM achieves significant improvements (t-test $p<0.05$) in most metrics over the suboptimal results. This highlights the advantages of integrating molecular knowledge through the multi-view graphs and the proposed trustworthy strategy.
IV-C Ablation studies
We conduct extensive ablation studies to evaluate the effectiveness of including the transcriptomics-derived RRI knowledge, the disease-specific R-RRI networks and the TFCP module. We accordingly remove the T-RRI network and three R-RRI networks, and replace the TFCP with TCP and a fully connected NN. The ablation experiments are performed on four tasks. Table III shows the results, from which we can observe that: 1) Both T-RRI and R-RRI individually improve prediction performance, demonstrating that transcriptomics data and imaging data each offer unique insights into disease pathology; 2) Integrating transcriptomics with imaging data provides the best prediction outcomes, showcasing that biomedical data from varied perspectives can deliver effective complementary insights for AD; 3) Compared to TCP, the proposed TFCP shows enhanced performance, and similarly, TCP outperforms NN, suggesting that combining certainty and uncertainty yields more robust confidence estimates.
TABLE IV: Top 15 imaging biomarkers identified for each task.
Task Modality Top 15 ROIsa
NC vs. AD VBM Amygdala_L, Hippocampus_R
FDG Cingulum_Post_L, ParaHippocampal_L, ParaHippocampal_R
AV45 Hippocampus_L, Hippocampus_R, ParaHippocampal_L, Cerebelum_3_L, Cerebelum_3_R, ParaHippocampal_R, Cerebelum_7b_R, Cingulum_Ant_R, Temporal_Pole_Sup_L, Cerebelum_10_L
NC vs. LMCI VBM Hippocampus_L, Amygdala_L, Hippocampus_R, Temporal_Mid_R
FDG Cingulum_Post_L, Cingulum_Post_R, ParaHippocampal_L, Angular_L, Angular_R, Thalamus_L, Occipital_Inf_L, Frontal_Med_Orb_L
AV45 ParaHippocampal_L, Hippocampus_R, Postcentral_L
NC vs. EMCI VBM Cerebelum_10_L, Hippocampus_L, Hippocampus_R, Amygdala_L, Cerebelum_9_L
FDG Cingulum_Post_L, ParaHippocampal_L, ParaHippocampal_R, Hippocampus_L, Angular_L, Angular_R, Cingulum_Post_R, Amygdala_L, Occipital_Mid_L
AV45 Cerebelum_3_L
EMCI vs. LMCI VBM Thalamus_R, Supp_Motor_Area_L, Hippocampus_L, Hippocampus_R
FDG Cingulum_Post_L, Putamen_R, Putamen_L, Cingulum_Post_R, Pallidum_L, Vermis_7, Pallidum_R, Temporal_Pole_Mid_R, Cerebelum_8_L, Vermis_8, Vermis_9
a: The biomarkers are ranked by combining all three modalities, and the top 15 features for each classification task are listed.
IV-D Evaluations of modality contributions
@@
To illustrate how multiple imaging modalities can offer complementary insights, we compared the classification performance using various combinations of imaging data sources on the NC vs. AD task. The results, displayed in Figure 2(a), reveal that: 1) different imaging modalities possess varying levels of discriminative capabilities, with VBM outperforming both FDG and AV45; 2) the integration of additional imaging modalities consistently yields better performance than using any subset alone, highlighting the unique contribution of each imaging modality and affirming the ability of our model in effectively modeling cross-modal information.
Figure 2: Performance comparison of (a) different modality combinations and (b) different hyperparameters for RRI thresholding on NC vs. AD task.
Figure 3: Coactivation maps of top ROIs obtained from NC vs. AD task.
Figure 4: Transcriptomics-based (top panel, grey edges) and radiomics-based (bottom panel, green edges) connectivity of top ROIs derived from each modality.
IV-E Hyperparameter analysis.
@@
We evaluate the hyperparameters for thresholding the T-RRI and R-RRI networks. The parameters $\lambda_{t}$ and $\lambda_{r}$ are automatically tuned from the NC vs. AD classification task within the range of $[0.1,0.2,\cdots,0.6]$. Figure 2(b) shows the grid search results, from which the optimal combination of $\lambda_{t}=0.2$ and $\lambda_{r}=0.1$ are selected in our experiments. Variations in the thresholding parameters have minimal impact on performance, demonstrating the robustness of RRI networks.
IV-F Identified biomarkers and functional analysis
Feature ablation experiments [45] are conducted to assess the importance of each imaging feature. Table IV lists the top 15 biomarkers identified for each task. The hippocampus ranks highly in both early and late prediction tasks, aligning well with its recognition as one of the first regions to be affected by AD [46, 47]. Among the top findings, several ROIs (e.g., left and right hippocampus, left posterior cingulum) are consistently present across various tasks, indicating their broad relevance. Some ROIs appear in most tasks (e.g., left amygdala, left parahippocampal gyrus), while some are specific to certain tasks (e.g., left thalamus in LMCI, right thalamus in EMCI vs. LMCI). This variability suggests that different regions may be involved at different AD stages.
We employ Neurosynth (www.neurosynth.org) to perform meta-analytic coactivation analysis, associating the identified ROIs with cognitive functions derived from 3,489 published neuroimaging studies. The coactivation map of top ROIs, as reported from NC vs. AD (refer to Table IV), is illustrated in Figure 3. The co-activated regions are primarily located in the subcortex (e.g., cingulate gyrus) and prefrontal lobe, which are known to be implicated in AD [48, 49].
We further visualize the connectivity of top ROIs underlying transcriptomics and each imaging modality. Specifically, we select the top 15 ROIs from each modality, extract their pairwise PCCs, and use the BrainNet Viewer (https://www.nitrc.org/projects/bnv/) for visualization. For each modality, distinct interaction patterns are observed between the transcriptomics-derived and radiomics-derived networks, illustrating how different biological levels offer unique insights from varying perspectives. This is especially marked in AV45, corresponding with the significant role of AV45 top ROIs (in NC vs. AD, 10 out of 15 ROIs are from the AV45) in classification. Further investigation is warranted into the identified ROIs and their role in AD progression.
@@ V Conclusion
In this paper, we propose TMM, a trustworthy enhanced multi-view multi-modal framework for AD prediction. We integrate upstream molecular knowledge with downstream radiomics information to effectively model relationships between brain ROIs and enhance sample representations. The proposed TFCP criterion successfully perceives sample-wise modal informativeness, facilitating dynamic fusion across multiple modalities. Our experimental results across various AD classification tasks validate the efficacy of the proposed TMM. The identified biomarkers align with existing research and demonstrate relevance to AD progression.

### v2

@@ III-C Modality-level Confidence: True-False-Harmonized Class Probability
In multimodal learning, data heterogeneity is a common challenge, with the discriminative capability of each modality varying across different samples. Therefore, it is crucial to develop algorithms that can adaptively perceive and respond to these variations in informativeness. By transforming the assessment of modal informativeness into an evaluation of the confidence levels associated with modal classification performance, we can effectively estimate the modality informativeness through its prediction confidence. Here, we propose a novel true-false-harmonized class probability (TFCP) to approximate the prediction confidence of each modality.
The confidence criterion measures how confident the model is in its predictions by correlating high prediction certainty with greater values and vice versa. For a classifier $f:\mathbf{X}\rightarrow\mathbf{y}$, the traditional confidence criterion utilizes the maximum class probability (MCP): $\text{MCP}(\mathbf{x}_{i})=P(\widehat{y_{i}}|\mathbf{x}_{i})$, where $\widehat{y_{i}}$ is the class with the largest softmax probability. However, it can be observed that this method assigns high confidence (i.e., high softmax probability) to both correct and incorrect predictions, resulting in overconfidence for failure predictions. True class probability (TCP) criterion [corbiere2019addressing] has been proposed to calibrate incorrect predictions. Instead of assigning the maximum class probability, TCP uses the softmax probability of the true class $y_{i}^{*}$ as prediction confidence: $\text{TCP}(\mathbf{x}_{i})=P(y_{i}^{*}|\mathbf{x}_{i})$, where $y_{i}^{*}$ is the class with the largest softmax probability. The TCP strategy exclusively models the certainty of the true class while neglecting the uncertainty of the untrue classes, which could result in a biased and unstable confidence approximation.
To this end, we propose considering both true and false class probabilities to aggregate evidence from these two perspectives. Specifically, we design a harmonized criterion, named $\text{Conf}_{\text{TFCP}}$, for the prediction confidence:
$\text{TFCP}(\mathbf{x}_{i})$ $=\frac{2}{1/{\text{TCP}(\mathbf{x}_{i})}+1/{1-\text{FCP}(\mathbf{x}_{i})}}$ (5)
@@
where $\text{FCP}(\mathbf{x}_{i})=P(\overline{y_{i}^{*}}|\mathbf{x}_{i})$ is the prediction probability of false (or untrue) class(es).
It should be noted that neither the TCP nor the FCP can be directly applied to test samples because they both require label information. To address this issue, we introduce two confidence networks to approximate the TCP and FCP, respectively. As shown in Figure 1(D), a TCP confidence network with parameters $\theta_{T}$ and an FCP confidence network with parameters $\theta_{F}$ are constructed based on the sample representations of each modality to generate certainty and uncertainty scores. These networks are trained to minimize the discrepancy between predicted and actual scores:
$\mathcal{L}_{\mathrm{Conf}}^{m}\!=\!\left\|\text{TFCP}(\mathbf{X}^{m})-\widehat{\text{TFCP}}(\mathbf{X}^{m},\theta^{m}_{T},\theta^{m}_{F})\right\|^{2}\!+\!\mathcal{L}_{\text{Cls}}^{m}.$ (6)
Here, both the TCP and FCP networks are built upon a classification network that is trained using the cross-entropy loss $\mathcal{L}_{\text{Cls}}^{m}$.
@@ IV-A Dataset
AHBA dataset. Brain-wide transcriptomics data are sourced from the Allen human brain atlas (AHBA, human.brain-map.org), including over 58k probes sampled across $3,702$ spatially distinct brain locations from six donors. We utilize the abagen toolbox [abagen] for data preprocessing and employ the AAL atlas to map brain locations to ROIs, resulting in a number of 15,633 genes expressed across 116 ROIs. To ensure the introduced prior knowledge closely aligns with the underlying pathogenic mechanisms of AD, we further filter the genes using large-scale AD meta-GWAS data from the International Genomics of Alzheimer’s Project (IGAP) [kunkle2019]. Specifically, we use the MAGMA tool to derive gene-level $p$-values from SNP-level GWAS results, keeping those with nominally significant $p$-value (i.e., $p<0.05$) into consideration. This process yields a total of 1,216 genes across 116 ROIs for constructing the T-RRI edge matrix.
@@
ADNI dataset. We conduct experimental validation using subjects from the Alzheimer’s Disease Neuroimaging Initiative database (ADNI, https://adni.loni.usc.edu/). Multimodal imaging data–including AV45 PET, FDG PET, and VBM MRI, along with corresponding diagnosis labels (NC, EMCI, LMCI, and AD) for each visit, are collected from 851 ADNI participants. Each imaging scan is processed according to established protocols [yao2019bioinfo], and measurements for 116 ROIs are derived from each modality using the AAL atlas. Detailed characteristic information and the number of subjects in each diagnostic group are listed in Table I.
Benchmark methods. We compare the proposed TMM with nine multimodal competitors, including three single-modal classifiers with early fusion (SVM, Lasso, XGBoost, and fully connected NN), three models with intermediate/late concatenation (GRidge [van2016better], BSPLSDA [singh2019diablo], and gated multimodal fusion (GMU [arevalo2017gated])), and two methods with advanced representation and fusion designs (Mogonet [2021MOGONET] and Dynamics [Han2022MultimodalDD]).
@@
Best is in Bold. Suboptimal results are underlined. ‘*’ indicates that our TMM is significantly better (t-test $p<0.05$) than the suboptimal method.
Evaluation metrics. We employ accuracy (ACC), F1 score (F1), and the area under the receiver operating characteristic curve (AUC) to evaluate the performance of the methods. Five-fold cross-validation is performed to calculate the mean and standard deviation of the results. T-tests are performed to evaluate the significance of improvements achieved by our method over state-of-the-art methods.
Implementation details. We develop the model using PyTorch 2.1.0 and Adam as the optimizer. Specifically, we train the model for 2000 epochs with a learning rate (lr) of 1e-3 and weight decay (wd) of 1e-4. We implement all experiments on an RTX 4090 GPU with 24GB of memory.
IV-B Comparison with the state-of-the-art
Four tasks, including NC vs. AD, NC vs. EMCI, NC vs. LMCI, and EMCI vs. LMCI, are designed to validate the performance of our proposed method. Since the transcriptomics knowledge is not specific to ADNI subjects and cannot be directly incorporated, we are limited to using only multimodal imaging data for the comparative methods.
Table II shows the comparison results where our model illustrates superior performance across four tasks. Particularly, TMM achieves significant improvements in most evaluation metrics over the suboptimal results. This highlights the advantages of integrating prior transcriptomics knowledge through the well-designed multi-view graphs and the proposed trustworthy strategy.
IV-C Ablation studies
@@
Effectiveness of T-RRI, R-RRI and TFCP. We conduct ablation studies to evaluate the effectiveness of including the transcriptomics-derived RRI knowledge and the TFCP module. We also examine the informativeness of disease-specific R-RRI networks and the efficacy of cross-attention in information consolidation.
For comparative analysis, we accordingly remove the T-RRI network and three R-RRI networks, replace the TFCP with TCP and a fully connected neural network (NN) and replace the cross-view and cross-modal attentions with NN. We conduct ablation experiments on xxx tasks. Table III shows the results, from which we have the following observations: 1).
Figure 2: Performance comparison of different modality combinations on NC vs. AD classification.
IV-D Evaluations of modality contributions and parameter analysis
@@
Modality contributions. To illustrate how multiple imaging modalities can offer complementary insights, we compared the classification performance using various combinations of imaging modalities. The results, displayed in Figure 2(b), reveal that: 1) different imaging modalities possess varying levels of discriminative capabilities, with VBM outperforming both FDG and AV45; 2) the integration of additional imaging modalities consistently yields better performance than using any subset alone, highlighting the unique contribution of each imaging modality and affirming the ability of our model in effectively modeling cross-modal information.
@@
Hyperparameter analysis. We evaluate the hyperparameters for thresholding in the T-RRI and R-RRI networks. The parameters $\lambda_{t}$ and $\lambda_{r}$ are automatically tuned from the NC vs. AD task within the range of $[0.1,0.2,\cdots,0.6]$. Figure 3(a) shows the grid search results, based on which we select the optimal parameter combination of $\lambda_{t}=0.2$ and $\lambda_{r}=0.1$ for use across our experiments. We observe that variations in the thresholding parameters have minimal impact on performance, demonstrating the robustness of our proposed method.
Figure 3: Performance comparison of difference hyperparameters.
IV-E Identified biomarkers and functional annotation
For each classification task, feature ablation experiments [Dai_2021_WACV] are conducted to assess and rank the importance of ROIs. Table LABEL:tab:biomarker lists the top xx ROIs identified from each task.
@@ V Conclusion
In this paper, we propose an adaptive multiomics integration framework introducing multi-view knowledge transfer for robust diagnosis classification. Graph contrastive pretraining is introduced to derive intra-omics patterns in an unsupervised manner. Cross-omics distillation facilitates the automatic transfer of inter-omics information from stronger to weaker modalities. Extensive experiments on benchmark datasets demonstrate the effectiveness of our approach. Concurrently, we have identified biomarkers that significantly impact disease inference, offering vital insights and pathways for elucidating pathological mechanisms and developing novel therapeutic approaches.